A water faucet in a house is h=8 m below the water surface in the local water tower. The radius of the orifice in the fully opened faucet is r=0.01 m. We assume non-viscous flow in this problem. How long does it take to fill a 20 liter bucket with water using this faucet????
Answer #1


speed of the water coming fro faucet = v =sqrt(2*g*h)


v = sqrt(2*9.8*8) = 12.52 m/s

volume flow rate = A*v

time required to fill 20 Liter = t = V/(A*v)

t = (20*10^-3)/(pi*0.01^2*12.52)


t = 5.1 s